The values of x which satisfies the equation (x+3)^4+(x+5)^4=16 are?

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
On expanding the above equation we get
2x4+32x3+204x2+608x+706=16
2x4+32x3+204x2+608x+690=0
2(x+3)(x+5)(x2+8x+23)=0
Appying quadratic equation formula we get the roots as
-3,-5,\small \sqrt{7}i-4,\small -\sqrt{7}i-4
Thanks and Regards
Shaik Aasif
askIITians faculty